OpenStudy (anonymous):

sum of the interior angles of an octagon? 180 360 1,080 or 3,640

OpenStudy (anonymous):

draw triangles in a standard octagon, depending on how much triangles u can fit, multiply that by 180.

OpenStudy (anonymous):

i meant which is the sum of the interior angles of an octagon?

jimthompson5910 (jim_thompson5910):

S = 180(n - 2) S = 180(8 - 2) S = 180(6) S = 1080 So the sum of the interior angles of an octagon is 1080 degrees